A 700-N man jumps out of a window into a fire net 10m below. The net stretches 2m before bringing the man to rest and tossing him back to the air. The maximum potential energy of the net, compared to its unstretched potential energy is?


1
Expert's answer
2021-03-24T19:42:25-0400

Let's apply the law of conservation of energy:


"PE_{man,i}+PE_{net,i}=PE_{man,f}+PE_{net,f},""W_{man}h_1+PE_{net,i}=-W_{man}h_2+PE_{net,f},""PE_{net,max}=PE_{net,f}-PE_{net,i}=W_{man}(h_1+h_2),""PE_{net,max}=700\ N\cdot(10\ m+2\ m)=8400\ J."